Jen Butler has been pricing​ Speed-Pass train fares for a group trip to New York. Three adults and four children must pay $ 113.  Two adults and three children must pay $ 80.  Find the price of the​ adult's ticket and the price of a​ child's ticket.
The price of a​ child's ticket is ​$

Answer by rothauserc(4718)

let x be the price of the adult's ticket and y be the price of the children's ticket
:
(1) 3x +4y = 113
:
(2) 2x +3y = 80
:
multiply equation 1 by 2 and equation 2 by 3
:
(3) 6x +8y = 226
:
(4) 6x +9y = 240
:
subtract equation 3 fro equation 4
:
y = 14
:
substitute y = 14 in equation 1
:
3x +4(14) = 113
:
3x +56 = 113
:
3x = 57
:
x = 19
:
*************************************************
adult's ticket costs $19 and children's costs $14
:
check the answer by substituting x=19 and y=14 in
equation 2
:
2(19) +3(14) = 80
:
38 +42 = 80
:
80 = 80
:
answer checks
**************************************************
: